Complete the ANOVA table. Round to two decimals if needed. Note that Factor = Between and Error = Within MS Source SS
Answer #1

Solution:

Given:

SSB = 53
SSW = 30

dfB = 3
dfW = 11

Part a)

MSB = ...........?

MSB = SSB / dfB

MSB = 53 / 3

MSB = 17.6667

MSB = 17.67

Part b)

MSW = ........?

MSW = SSW / dfW

MSW = 30 / 11

MSW = 2.7273

MSW = 2.73

Part c)

F = ..........?

F = MSB / MSW

F = 17.67 / 2.73

F = 6.47